celticcat:

I'm confused!! Just seen an answer on another website. Simplify 2 + 3 - 4 (5*4). Someone answered it as 20 and the questioner said it was right! Is this some new weird algebra or what!!???

Vocaloid:

pay no mind. it's just people who don't understand PEMDAS/order of operations.

Vocaloid:

you can only get 20 if you go from left to right and ignore order of operations 2+3 = 5 5 - 4 = 1 1*(5*4) = 20 which is incorrect

celticcat:

Yea i thought so. When I tried to explain that they said they had used pemdas lol!!

celticcat:

Then they said that's the way we do it in Kansas! lol!